OpenStudy (anonymous):

How do you solve the inequality, x squared + 3x - 4 is less than or equal to 0 ?

OpenStudy (anonymous):

\[x^{2} +3x -4 \le 0\] forget the inequality and solve like this: \[x^{2} +3x -4 = 0\] When you get the values for X then X must be less than or equal to those values

OpenStudy (anonymous):

well actually i was wrong. X must be inside those values so let's say the answer is -4 and 1 then \[-4 \le x \le 1\]

OpenStudy (anonymous):

thanks :)